OpenStudy (anonymous):

2x-y=5 9x-4y=23 use method of substitution @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

2x-y=5 -2x -2x -y=-2x+5 y=2x-5

HanAkoSolo (jamierox4ev3r):

9x-4(2x-5)=23 9x-8x+20=23 x+20=23 -20 -20 x=3

OpenStudy (alexwee123):

you just asked a problem like this @flower93 give the problem a try

HanAkoSolo (jamierox4ev3r):

2(3)-y=5 6-y=5 -6 -6 -y=-1 y=1

HanAkoSolo (jamierox4ev3r):

9(3)-4y=23 27-4y=23 -27 -27 -4y=-4 y=-1 I wasn't giving answers, i was showing my work then showing the solution i arrived at @mth3v4

HanAkoSolo (jamierox4ev3r):

lol okay but i don't think she knew where to begin... :P sorry, normally im a stickler for not giving out answers i must be tired or something O.o

OpenStudy (anonymous):

Exactly i didnt kn where to begin and @mth3v4 wasnt begin very helpful or showing me how.

OpenStudy (anonymous):

............

OpenStudy (anonymous):

Thank u @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

lol np @flower93 sorry @mth3v4

OpenStudy (anonymous):

lol

HanAkoSolo (jamierox4ev3r):

:P

OpenStudy (anonymous):

you always gave up flower

OpenStudy (anonymous):

you hardly answer any of my questions

HanAkoSolo (jamierox4ev3r):

hmmmm says the guy who says to give encouragement O.O lol jk

OpenStudy (anonymous):

Anyways thanks for your help, i appreciate you trying to help me out @mth3v4

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i had alot of ppl doing worse

OpenStudy (anonymous):

actually cursing me and i also did not help :P

HanAkoSolo (jamierox4ev3r):

lol who would help at that point.

OpenStudy (anonymous):

well they were just waiting for answers but that was before OS got a new makeover

HanAkoSolo (jamierox4ev3r):

ah i see how long have u been on here anyhow?

OpenStudy (anonymous):

almost longer than many ppl have been here

HanAkoSolo (jamierox4ev3r):

except for satellite and preetha and the other legendary users

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i been here longer than some mods :P

OpenStudy (anonymous):

and almost all ambies

HanAkoSolo (jamierox4ev3r):

duude, pretty awesome . I'm applying myself to become an ambie.

OpenStudy (anonymous):

y

HanAkoSolo (jamierox4ev3r):

idk, ithink it would be a good experience and i love this site and would like to make it better in my own way, so i just have to wait and see....

OpenStudy (anonymous):

just to say .... if you are thinking it will help you to become a mod ....no it wont